# Optimization-Free Fast Optimal Control: Bang-Ride Property, Monotonicity, and Applications to Fast Battery Charging

### v1

@@ 3.1 Introduction and System Properties
This subsection briefly recalls the system property for achieving bang-bang control [12], which is also useful for bang-ride control. The main difference in establishing bang-ride control from bang-bang control lies in guaranteeing the non-zero $\lambda$, which requires addressing the additional path constraints. This is presented in the next subsection.
@@
Assumption 3.2 is related to controllability. For linear systems, $M_{n}$ is the controllability matrix and thus Assumption 3.2 implies linear controllability. For nonlinear systems, Assumption 3.2 implies locally strong accessibility [19], i.e., one version of local controllability. It also implies that the system is feedback linearizable [19] and differentially flat [25]. More detailed discussions are presented in Appendix A.
@@
Lemma 2 is a variant of [12, Thm. 1] and [10, Thm. 6] following the same proof. The differences are that no constraint and no state constraint are considered in [12] and [10], respectively. These simplified problems lead to $\lambda\not=0$ straightforwardly. In contrast, due to state constraints in this work, $\lambda\not=0$ is explicitly assumed in Lemma 2 and needs to be achieved through complex regularity conditions, as done in the next subsection.
@@ 3.2 Qualification and Switching of Constraints
We investigate conditions for achieving a non-zero $\lambda$, as required in Lemma 2. We exploit a series of regularity conditions, including constraint qualifications and conditions on constraint switching. Constraint qualifications are typically rank conditions on the constraints. Regularity on constraint switching excludes potential discontinuities of $\lambda$ caused by the activation/deactivation of state constraints. Some regularity conditions in this subsection are from the classical optimal control literature, but have not been used to study the bang-ride property.
@@
When $t_{f}$ is free, one common choice is $l\equiv 1$, leading to the time-optimal control problem. Assumption 3.3b generalizes from $l\equiv 1$ to a non-zero objective function. Assumption 3.3a and c extend [12] to incorporate the terminal constraint and a non-zero objective, respectively. [^7] Equation (17) is satisfied vacuously if $z$ is absent.
When there is no path constraint, Assumption 3.3 achieves non-zero $\lambda$, analogous to the classical result [22]. In this work, we need to additionally address the state constraints, which lead to discontinuities of the costate $\lambda$. Following [13, Chap.  5.4], first define the index set of active state constraints right after and right before time $t$:
@@ 3.3 Main Result for Bang-Ride Control
Theorem 4 shows that combining the system property, related to controllability, and the regularity of constraint switching leads to bang-ride control. These conditions are sufficient. As shown in Sections 4 and 5, bang-ride control can also be achieved by monotone systems that are not controllable. However, the regularity of constraint switching remains important throughout this work.
@@ 3.4 System Decomposition
Even if the complete system may not satisfy the system properties in Assumption 3.2, an important situation, pointed out in [10], is when only a subpart of the system affects the optimal control solution and satisfies the properties. One straightforward situation is when the system is in the form:
@@
and all the functions in problem (2) depend only on $x_{1}$. Then, besides the regularity conditions on the constraints, the system properties in Assumption 3.2 only need to be satisfied by (18a) for achieving the bang-ride property.
A more interesting situation occurs when the system is not in the form of (18) but can be transformed to this form via a coordinate transformation. We illustrate this via linear systems. Suppose that problem (2) has the form:
@@
As only $y$ affects the optimal solution, solving problem (19) using system model (21) leads to the same optimal solution. Therefore, to ensure the bang-ride property, only the subsystem $\dot{\bar{x}}_{1}=\bar{A}_{11}\bar{x}_{1}+\bar{B}_{1}u$ needs to satisfy the system properties in Assumption 3.2, i.e., being controllable.
@@ 4.1 Introduction
The positive $\lambda^{\top}g^{\star}$ in (22) leads to a special class of bang-ride control, compared to being non-zero in (14) for a general bang-ride control. When there is only an input constraint $u\in[u_{\mathrm{min}},u_{\mathrm{max}}]$, (22) and (23) lead to $u^{\star}=u_{\mathrm{max}}$. When other path constraints are present, (23) shows that the resulting input is the maximum feasible input. Our goal is to develop conditions for (22).
@@ 4.2 Limitation of Controllability
Nevertheless, both systems are controllable in Example 2.1, and thus controllability is not sufficiently informative to distinguish (22) from general bang-ride control. This motivates us to exploit the monotonicity of the optimal control problem. As shown later in Example 5.1, this direction leads to system properties that can distinguish the two systems in Example 2.1.
@@ 4.3 Monotonicity
Assumption 4.1 requires the terminal components to be monotonically decreasing. Note that the inactive terminal constraints, whose associated multipliers are zero based on (9b), are excluded from the above restriction via set $\mathcal{I}_{f}$.
Minimizing a decreasing terminal objective $\phi$ leads to an incentive for a larger state, which can be achieved by applying a larger input for special system classes. This is related to the monotonicity of the system.
@@
Assumption 4.2a is useful for establishing (22). Moreover, based on [15, Prop. 3.3], Assumption 4.2 indicates that the system is a monotone system [15] in a neighborhood of the optimal trajectory, i.e., a larger initial state and larger input lead to a larger state trajectory. Assumption 4.2b can also be guaranteed by the following more explicit conditions:
@@
Finally, the path constraints and the path objective $l$ should also behave well so that the optimal solution prefers a larger input.
@@
Assumption 4.3a requires $l$ to be monotonically decreasing, analogous to Assumption 4.1, and thus a larger state is desired to decrease the objective. Assumption 4.3b imposes that changing $u$ and $x$ have opposite effects on the constraint $\bar{s}_{i}$.
We can now obtain an intuition about how the properties in Assumptions 4.1, 4.2, and 4.3 affect the optimal solution. The monotonicity of the terminal components and stage cost shows that an optimal solution prefers a larger state. This leads to a preference for a larger input, due to the monotonicity of the system. Finally, the larger input and state do not tend to violate the constraints, as they have the opposite effect on the constraint $\bar{s}_{i}$. Under these properties, the optimal input is likely to be the maximum feasible input.
@@
Theorem 2 shows that the optimal input is the maximum feasible input under the monotonicity assumptions. An important special case is when there is only an input bound, as in the bang-bang control, then $u^{\star}=u_{\mathrm{max}}$. Assumptions 4.1, 4.2, and 4.3 are analogous to the properties in [6]. In this work, we take a different analytical approach by exploiting PMP. This allows us to address the regularity conditions formally. More importantly, it opens up new ways to generalize the result in Remark 4 and Section 5.
@@
For a linear system (20), Assumption 4.2 holds iff $B>0$ and $A$ is a Metzler matrix. Moreover, in (19), the properties in the assumptions may be affected by a similarity transformation of the linear system (20). For example, a linear system with $B<0$ can be transformed into an equivalent model with $B>0$. The challenge is to find a transformation to satisfy all assumptions simultaneously. Developing properties invariant to similarity transformation is desired, which is beyond the scope of this article.
@@ 4.4 Connection to Hybrid Simulation
Proposition 6 shows that the solution of the hybrid simulation satisfies (23) for its state trajectory. Recall from Theorem 2 that, under its assumptions, (23) becomes a necessary optimality condition. Therefore, if the solution of the hybrid simulation satisfies the assumptions in Theorem 2, the solution also satisfies the necessary optimality condition (23). In this case, the solution of the hybrid simulation is likely to be optimal.
@@ 5.1 Introduction and General Results
Some physical constraints do not satisfy $\partial s/\partial x\leqslant 0$ in (24). An effort is made in [6] to address this limitation; however, their result is limited to linear constraints. Moreover, instead of requiring $\lambda>0$, $\lambda^{\top}(t)g$ can remain positive even if some elements of $\lambda(t)$ become negative. This phenomenon is related to the problem of external positivity:
@@
The above generality of Theorem 3 is achieved by the additional restrictions on the system in Assumption 5.1b and on the relative sensitivity of the constraints in Assumption 5.1c. Nonetheless, we view Theorem 3 as an important foundation for a more general solution of the problem in future research.
@@
Recall the optimal control problem (12), whose system is in the system class (26), and the solutions of the hybrid simulation in Fig. 1 for $a_{1}=0$, $a_{2}=-1$ and in Fig. 2 for $a_{1}=-1$, $a_{2}=0$. From its terminal objective $\phi(t_{f})=-x_{1}(t_{f})$ and (9a), we have $\lambda(t_{f})=[1\text{ }0]^{\top}$, showing $\lambda_{1}(t_{f})=1>\lambda_{2}(t_{f})=0$. Then, the case with $a_{1}=0$ and $a_{2}=-1$ satisfies Assumptions 5.1ab due to $a_{1}>a_{2}$. However, the case with $a_{1}=-1$ and $a_{2}=0$ does not satisfy these assumptions.
@@ 5.2 Special Setting for Application
6 Application: Fast Battery Charging
6.1 Battery Charging with the Single-Particle Model (SPM)
Here, the theory is illustrated by applying it to a battery fast-charging problem using the SPM [20]. The SPM describes the evolution of lithium concentration $c_{j}(r,t)$ over a spatial coordinate $r\in[0,R_{j}]$:
@@ 6.1 Battery Charging with the Single-Particle Model (SPM)
where $D_{j,s}$ is a positive parameter, and $j\in\{+,-\}$ denotes the index for the positive or negative electrode. The boundary conditions are
@@
For our purpose, only the surface concentration $c_{j,\mathrm{s}}(t):=c_{j}(R_{j},t)$ and the average concentration $c_{j,\mathrm{ave}}$ [18] are of interest. Regarding these two quantities as two outputs leads to an approximation of (29) in the form of a linear state-space model [^11] While each electrode is approximated by a $3$-dimensional linear system, leading to $6$ states in total, both electrodes contain an integrator, and thus one redundant state for integration is eliminated. [5, 18]:
@@ 6.2 Analysis of the Fast-Charging Problem
It is well-known from the numerical results [1] that the optimal input is the CC-CV profile, which is consistent with the hybrid simulation, as also shown in Fig. 4. We provide a novel theoretical justification for this using Proposition 4.
@@
Recall that model (30) is already in the form of (26) and is thus a monotone system. For Assumption 5.2a, $\mathrm{SOC}$ depends on $c_{+,\mathrm{ave}}$ and thus only on the state $x_{1}$ due to (31). We can compute $\lambda(t_{f}^{-})$ from (9a) and thus from $\partial\mathrm{SOC}/\partial x$. First, SOC satisfies
@@
In summary, Assumption 5.2 is satisfied, and thus $u=\max\mathcal{D}(x)$ becomes a necessary optimality condition, based on Proposition 4. Then, we can conclude that the hybrid simulation and the CC-CV profile satisfy this necessary optimality condition, based on Proposition 6. Although its optimality is well-known empirically, the above formal theoretical analysis is novel.
Note that the bang-ride behavior can also be explained by Theorem 4 as system (30) is controllable. However, Theorem 2 and the result in [6] are not applicable, as they require the nonlinear constraint to be decreasing in state. The analysis in [5] is limited to linear constraints and thus is not applicable here.
@@ 7 CONCLUSIONS
Motivated by the properties of the heuristic battery charging profiles and the hybrid simulation, this work exploits PMP to develop theoretical conditions under which the optimal input has these properties. The theory shows that optimal bang-ride control can be achieved by (nonlinear local) controllability and regulated constraint switching. Moreover, a special class of bang-ride control, which applies the maximum feasible input, can be achieved by the monotonicity of the system dynamics, constraints, and objective. In settings with more general nonlinear constraints and linear systems, the above bang-ride control can also be established by the external positivity of costate dynamics and restrictions on the relative sensitivity of the constraints. In these settings, we also justify that the heuristic input profiles and hybrid simulation satisfy the necessary optimality condition.
An important direction for future work is addressing more general constraints in the optimal control problem. For example, some constraints are increasing while others are decreasing functions of the state, or a constraint is increasing in some states and decreasing in the other states. These constraints lead to more complex behavior for the costates. While these cases are considered in Theorem 3, its assumptions do not cover all practical situations. The above generalization relates to the external positivity of the costate dynamics, as formulated in Problem 1. However, the complexity of external positivity analysis suggests that theoretical results, at least the sufficient conditions for maximum input being optimal, may need to be tailored to specific problem settings, as done in Section 5.2. Another direction is considering broader system classes, such as partial differential equations or nonlinear systems with input nonlinearity.
@@ 7.1 Discussion: Nonlinear Controllability, Feedback Linearization, and Differential Flatness
This property means that, from every state, the system can reach an open region close to the starting point, indicating a certain level of capability of being controlled. Note that this is a local concept.
Locally strong accessibility can be determined by the rank (dimension) test of a nonlinear generalization of the controllability matrix (controllable space), called strong accessibility algebra, whose formal definition can be found in [19]. Then it can be proved that Assumption 3.2 implies locally strong accessibility.
@@
Based on [19, Prop. 3.20], every element of the strong accessibility algebra has the form
@@
Directly follows from [19, Cor. 6.17] and Assumption 3.2.

### v3

@@ 3.1 Introduction and System Properties
This subsection briefly recalls the system property to achieve bang-bang control [13], which is also useful for bang-ride control. The main difference in establishing bang-ride control from bang-bang control lies in guaranteeing non-zero $\lambda$, which requires addressing the additional path constraints. This is presented in the next subsection.
@@
Assumption 3.2 is related to controllability. For linear systems, $M_{n}$ is the controllability matrix, and thus Assumption 3.2 implies linear controllability. For nonlinear systems, Assumption 3.2 implies locally strong accessibility [20], i.e., one version of local controllability. It also implies that the system is feedback linearizable [20] and differentially flat [26]. More detailed discussions are presented in Appendix A.
@@
Lemma 2 is a variant of [13, Thm. 1] and [11, Thm. 6] following the same proof. The differences are that no constraints are considered and no state constraints are in [13] and [11], respectively. These simplified problems lead to $\lambda\not=0$ in a straightforward way. In contrast, due to state constraints in this work, $\lambda\not=0$ is explicitly assumed in Lemma 2 and needs to be achieved through additional regularity conditions, as done in the next subsection.
@@ 3.2 Qualification and Switching of Constraints
We investigate conditions for achieving a non-zero $\lambda$, as required in Lemma 2. We exploit a series of regularity conditions, including constraint qualifications and conditions on constraint switching. Constraint qualifications are typically rank conditions on the constraints. Regularity on constraint switching excludes potential discontinuities of $\lambda$ caused by the activation/deactivation of state constraints. Some regularity conditions in this subsection are classical, but have not been used to study the bang-ride property.
@@
When $t_{f}$ is free, a common choice is $l\equiv 1$, leading to the time-optimal control problem. Assumption 3.3b generalizes from $l\equiv 1$ to a non-zero objective function. Assumption 3.3a and c extend [13] to incorporate the terminal constraint and a non-zero objective, respectively. [^7] Equation (17) is satisfied vacuously if $z$ is absent.
When there is no path constraint, Assumption 3.3 achieves non-zero $\lambda$, analogous to the classical result [23]. In this work, potential discontinuities of $\lambda$ caused by state constraints need to be addressed. We exploit classical regularity conditions in [14, Chap.  5.4] for our purposes. First, define the index set of active state constraints right after and right before time $t$:
@@ 3.3 Main Result for Bang-Ride Control
Theorem 4 shows that combining the system property, related to controllability, and the regularity of constraint switching leads to bang-ride control. These conditions are sufficient. As shown in Sections 4 and 5, bang-ride control can also be achieved by monotone systems that are not controllable. However, the regularity of constraint switching remains important.
@@ 3.4 System Decomposition
Even if the complete system may not satisfy Assumption 3.2, an important situation, pointed out in [11], is when only a subpart of the system affects the optimal control solution and satisfies the assumption. One straightforward situation is when the system is in the form:
@@
and all the functions in problem (2) depend only on $x_{1}$. Then, in addition to the regularity conditions in the constraints, the system properties in Assumption 3.2 only need to be satisfied by (18a) to achieve the bang-ride property.
A more interesting situation occurs when the system is not in the form of (18) but can be transformed to this form via a coordinate transformation. We illustrate this considering linear systems. Suppose that problem (2) has the form:
@@
As only $y$ affects the optimal solution, solving problem (19) using system model (21) leads to the same optimal solution. Therefore, to ensure the bang-ride property, only the subsystem $\dot{\bar{x}}_{1}=\bar{A}_{11}\bar{x}_{1}+\bar{B}_{1}u$ needs to satisfy Assumption 3.2, i.e., being controllable.
@@ 4.1 Introduction
The positive $\lambda^{\top}g^{\star}$ in (22) leads to a special class of bang-ride control, compared to being non-zero in (14). When there is only an input constraint $u\in[u_{\mathrm{min}},u_{\mathrm{max}}]$, (22) and (23) lead to $u^{\star}=u_{\mathrm{max}}$. When other path constraints are present, (23) shows that the optimal input is the maximum feasible input. Our goal is to develop conditions for (22).
@@ 4.2 Limitation of Controllability
Nevertheless, as both systems in Example 2.1 are controllable, controllability is not informative enough to distinguish (22) from the general bang-ride control. This motivates us to exploit monotonicity. As shown later in Example 5.1, this direction leads to system properties that can distinguish the two systems in Example 2.1.
@@ 4.3 Monotonicity
Assumption 4.1 requires the terminal components to decrease monotonically. Note that the inactive terminal constraints, whose associated multipliers are zero based on (9b), are excluded from the above restriction via set $\mathcal{I}_{f}$.
Minimizing a decreasing terminal objective $\phi$ leads to an incentive for a larger state, which can be achieved by applying a larger input for special system classes. This is related to monotonicity of the system.
@@
Assumption 4.2a is useful for establishing (22). Moreover, based on [16, Prop. 3.3], Assumption 4.2 indicates that the system is a monotone system [16] in a neighborhood of the optimal trajectory, i.e., a larger initial state and larger input lead to a larger state trajectory. Assumption 4.2b can also be guaranteed by the following more explicit conditions:
@@
Finally, the path constraints and path objective $l$ should behave well so that the optimal solution prefers a larger input.
@@
Assumption 4.3a requires $l$ to be monotonically decreasing, analogously to Assumption 4.1, and thus a larger state is desired to decrease the objective. Assumption 4.3b imposes that changing $u$ and $x$ have opposite effects on constraint $\bar{s}_{i}$.
We can now obtain an intuition about how the properties in Assumptions 4.1, 4.2, and 4.3 affect the optimal solution. The monotonicity of the terminal components and the stage cost shows that an optimal solution prefers a larger state. This leads to a preference for a larger input, due to the monotonicity of the system. Finally, the larger input and the larger state do not tend to violate the constraints, as they have the opposite effect on the constraint $\bar{s}_{i}$. Under these properties, the optimal input is likely to be the maximum feasible input.
@@
Theorem 2 shows that the optimal input is the maximum feasible input under the monotonicity assumptions. An important special case is when there is only an input bound, as in the bang-bang control, then $u^{\star}=u_{\mathrm{max}}$. Assumptions 4.1, 4.2, and 4.3 are analogous to the properties in [7, 10]. In this work, we take a different analytical approach by exploiting PMP. This allows us to address the regularity conditions formally and make an explicit connection to the bang-ride control. More importantly, it opens up new ways to generalize the result in Remark 4 and Section 5.
@@
For a linear system (20), Assumption 4.2 holds iff $B>0$ and $A$ is a Metzler matrix. Moreover, in (19), the properties in the assumptions may be affected by a similarity transformation. For example, a linear system with $B<0$ can be transformed into an equivalent model with $B>0$. The challenge is to find a transformation to satisfy all assumptions simultaneously. Developing properties invariant to similarity transformation is desired, which is beyond the scope of this article.
@@ 4.4 Connection to Hybrid Simulation
Proposition 6 shows that the solution of the hybrid simulation satisfies (23) for its state trajectory. Recall that, under the assumptions of Theorem 2, (23) becomes a necessary optimality condition. Therefore, if these assumptions are satisfied by the solution, the solution also satisfies the necessary optimality condition (23). In this case, the solution of the hybrid simulation is likely to be optimal.
@@ 5.1 Introduction and General Results
Some physical constraints do not satisfy the condition $\partial s/\partial x\leqslant 0$ in (24). An effort is made in [7] to address this limitation; however, their result is limited to linear constraints. Moreover, instead of requiring $\lambda>0$, $\lambda^{\top}(t)g$ can remain positive even if some elements of $\lambda(t)$ become negative. We establish a novel connection between this phenomenon and the problem of external positivity:
@@
The above generality of Theorem 3 is achieved by the additional restrictions on the system in Assumption 5.1b and on the relative sensitivity of the constraints in Assumption 5.1c. Nonetheless, we view Theorem 3 as an important foundation for a more general solution in future research.
@@
Recall problem (12), whose system is in the system class (26), and the solutions of the hybrid simulation in Fig. 1 for $a_{1}=0$, $a_{2}=-1$ and in Fig. 2 for $a_{1}=-1$, $a_{2}=0$. From its terminal objective $\phi(t_{f})=-x_{1}(t_{f})$ and (9a), we have $\lambda(t_{f})=[1\text{ }0]^{\top}$, showing $\lambda_{1}(t_{f})=1>\lambda_{2}(t_{f})=0$. Then, the case with $a_{1}=0$ and $a_{2}=-1$ satisfies Assumptions 5.1ab due to $a_{1}>a_{2}$. However, the case with $a_{1}=-1$ and $a_{2}=0$ does not satisfy these assumptions.
@@ 5.2 Special Setting for Application
6 Application Example: Fast Battery Charging
6.1 Charging with the Single-Particle Model (SPM)
The theory is illustrated by applying it to a battery fast-charging problem using the SPM [21]. The SPM describes the evolution of lithium concentration $c_{j}(r,t)$ over a spatial coordinate $r\in[0,R_{j}]$:
@@ 6.1 Charging with the Single-Particle Model (SPM)
where $D_{j,s}$ is a positive parameter and $j\in\{+,-\}$ denotes the index of the positive or negative electrode. The boundary conditions are
@@
For our purposes, only the surface concentration $c_{j,\mathrm{s}}(t):=c_{j}(R_{j},t)$ and the average concentration $c_{j,\mathrm{ave}}$ [19] are of interest. Regarding these two quantities as two outputs leads to an approximation of (29) in the form of a linear state-space model [^11] While each electrode is approximated by a $3$-dimensional linear system, leading to $6$ states in total, both electrodes contain an integrator, and thus one redundant state for integration is eliminated. [6, 19]:
@@ 6.2 Analysis of the Fast-Charging Problem
It is well-known from numerical results [1] that the optimal input is the CC-CV profile, which is consistent with the hybrid simulation, as also shown in Fig. 4. We provide a novel theoretical justification for this using Proposition 4.
@@
Recall that model (30) is in the form of (26) and is a monotone system. For Assumption 5.2a, $\mathrm{SOC}$ depends on $c_{+,\mathrm{ave}}$ and thus only on the state $x_{1}$ due to (31). We can compute $\lambda(t_{f}^{-})$ from (9a) and thus from $\partial\mathrm{SOC}/\partial x$. First, SOC satisfies
@@
In summary, Assumption 5.2 is satisfied, and thus $u=\max\mathcal{D}(x)$ becomes a necessary optimality condition, based on Proposition 4. Then, the hybrid simulation and the CC-CV profile satisfy this necessary optimality condition, based on Proposition 6. Although its optimality is well-known empirically, the above formal theoretical analysis is novel.
The bang-ride behavior can also be explained by Theorem 4 as (30) is controllable. However, Theorem 2 and the result in [7] are not applicable, as they require the nonlinear constraint to be decreasing in state. The analysis in [6] is limited to linear constraints and thus is not applicable here.
@@ 7 CONCLUSIONS
This work has provided theoretical conditions under which heuristic input profiles and hybrid simulation methods in time-critical control problems coincide with necessary optimality. Using Pontryagin’s maximum principle, we showed that bang-ride structures arise under controllability and regular constraint switching, and that maximum-feasible inputs become necessary under monotonicity or external positivity of the costate dynamics. These results justify widely used charging heuristics and provide efficient, optimization-free certificates for their optimality.
Future research should address settings with mixed monotonicity, where some constraints increase and others decrease with the state, as well as system classes beyond finite-dimensional linear dynamics. Such generalizations may require tailored notions of external positivity or problem-specific sufficient conditions, and could extend the applicability of these results to PDE systems or nonlinear dynamics with input nonlinearities.
@@ 7.1 Discussion: Nonlinear Controllability, Feedback Linearization, and Differential Flatness
This property means that, from every state, the system can reach an open region close to the starting point, indicating a capability of being controlled. Note that this is a local concept.
Locally strong accessibility can be determined by the rank (dimension) test of a nonlinear generalization of the controllability matrix (controllable space), called strong accessibility algebra, whose formal definition is in [20]. Then Assumption 3.2 implies locally strong accessibility.
@@
Based on [20, Prop. 3.20], every element of the strong accessibility algebra has the form
@@
Directly follows from [20, Cor. 6.17] and Assumption 3.2.